A tennis ball is thrown with an initial upward velocity of 64 feet per second. How long will it take for the ball to hit the ground? Use the formula in which the height above ground after t seconds is given by the quadratic function h = vt – 16t2, where h represents the height of the object in feet, and v represents the object's initial velocity in feet per second.
The answer is supposed to be in seconds, but I am not at all sure how to get there.

Answer by lwsshak3(11628)

A tennis ball is thrown with an initial upward velocity of 64 feet per second. How long will it take for the ball to hit the ground? Use the formula in which the height above ground after t seconds is given by the quadratic function h = vt – 16t^2, where h represents the height of the object in feet, and v represents the object's initial velocity in feet per second.
The answer is supposed to be in seconds, but I am not at all sure how to get there.
**
When the ball hits the ground,
h=0
vt–16t^2=0
64t-16t^2=0
..
t(64-16t)=0
t=0 (reject, not reasonable)
or
64-16t=0
16t=64
t=64/16=4
How long will it take for the ball to hit the ground? 4 seconds